# Two-phase formation of galaxies: the coevolution between galaxies and dark matter halos

### v2

@@ 2.1.1 Stellar kinematics and decomposition
Here, $f_{1}$ and $f_{2}$ are the initial and final values of $f_{\rm thin}$, respectively, introduced to describe the shape of $f_{\rm thin}$ seen in Fig. 2(c) at $t<t_{1}$ and $t\geqslant t_{2}$. The change of $f_{\rm thin}$ in the transitional interval between $t_{1}$ and $t_{2}$ is described by a linear function. The fitting result for m12f is shown by the blue dashed curve in Fig. 2(c), with the duration of the transitional phase ($z\approx 0.5$11, or $t\approx 5.5$$8.5\,{\rm Gyr}$) indicated by the grey shade. To avoid ambiguity, we will denote $t_{1}$ and $t_{2}$ as $t_{\rm thin}^{\rm(start)}$ and $t_{\rm thin}^{\rm(end)}$, respectively. We use the mean value of the two,
@@ 2.2 Dark matter halos hosting the galaxies
The exact formation redshift is evaluated by linearly interpolating the redshift between the two adjacent snapshots where $M_{\rm vir}$ first crosses the half of its final mass. This time has been commonly used in literature to characterize the assembly of dark matter halos (Gao et al., 2005; Li et al., 2008), and was found to correlate with $t_{\gamma}$ (see Appendix B of MCW24). Another related formation time, which is also commonly used in the literature to characterize halo assembly, is $t_{\rm h,Vpeak}$, defined as the cosmic time when the maximum circular velocity ($V_{\rm max}$) of the halo reaches its peak along the main-branch history. This time was found to be sensitive to violent changes in the halo mass, and has been used to incorporate effects of recent major mergers (Behroozi et al., 2014).
In Table 2, we list the properties of galaxies/halos defined in this subsection and the last. Using these definitions, we obtain the characteristic times for each of the galaxies/halos in our sample, and we list them in Table 1. In general, the piecewise functions introduced in §2.1 fit well the evolution of the properties for most galaxies in our sample. The fitting is less satisfactory for m12w, m12r and m12z. Galaxies m12w and m12z have never reached to the disk-dominated stage, so that the intervals $_{\rm thin}^{\rm(start)}, tt$$_{\rm thin}^{\rm(end)}$] and $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(start)}$, $t_{\rm\scriptscriptstyle\mathcal{H}}^{\rm(end)}$] are wide, and the corresponding transition times, $t_{\rm thin}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}$, are not well constrained. Galaxy m12r has a massive satellite at a distance of $390\,{\rm{kpc}}$ at $z=0$, with stellar mass of $M_{*}\approx 2.8\times 10^{9}{\rm M_{\odot}}$ and a recent pericentric passage at $30\,{\rm{kpc}}$ only about $0.7\,{\rm Gyr}$ ago ($z=0.05$; Samuel et al. 2020). Consequently, the disk of m12r is strongly perturbed, which causes significant fluctuations in $f_{\rm thin}$ and $\mathcal{H}$. The disk of m12r is still preserved during the interaction, primarily because of the largely prograde motion of the satellite with respect to the disk. The exceptional dynamics of these three galaxies also have impact on their burstiness. Galaxies m12w and m12r show significant fluctuations in $\mathcal{B}$ around $0.2$ (the threshold to define $t_{\rm\scriptscriptstyle\mathcal{B}}$), while galaxy m12z never reaches this threshold and so its $t_{\rm\scriptscriptstyle\mathcal{B}}$ is absent in Table 1. Since transitions are quite vague for the three galaxies, we analyze and discuss them separately in the following.
@@ 3 Co-evolution pattern between galaxies and halos
Figure 3: Correlation between halo transition time ($t_{\gamma}$) and galaxy transition times. Each panel in the upper row uses one definition of galaxy transition time: (A), $t_{\rm thin}$; (B), $t_{\rm\scriptscriptstyle\mathcal{H}}$; (C), $t_{\rm\scriptscriptstyle\mathcal{B}}$; (D), $t_{\rm bulge,1/2}$. Each marker represents one galaxy in our sample, with error bar spanning from the start to the end of the transition interval (for $_{\gamma}, tt$$_{\rm thin}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}$) or from 20% to 80% formation time of bulge mass (for $t_{\rm bulge,1/2}$). Markers with grey edges represent the peculiar galaxies (m12w, m12r and m12z) excluded from correlation analysis. For each pair of variables, the Spearman correlation coefficient ($\rho$) is indicated in the corresponding panel (see also Table 3 for a full list of $\rho$ and $p$-values), and black line shows a linear fitting. Each lower panel shows $\Delta T_{\rm dyn}$, the duration of the transition interval corresponding to the upper panel, normalized by the dynamical time of the host halo. This figure demonstrates the strong correlation between halo assembly and galaxy assembly, and implies that the transitions of galaxies are driven by the transition of halo assembly. See §3 for details.
@@
Besides the strong correlations between galaxy properties and halo assembly, there are significant variations among individual galaxies, indicating that processes other than the halo accretion may also play important roles in determining the transitions in galaxy properties. For example, the galaxy m12m has a $t_{\rm thin}$ about $2\,{\rm Gyr}$ earlier than that expected from the linear fit. A careful examination reveals that m12m has a massive satellite whose orbital motion aligns with the disk rotation of m12m at $z\approx 3$. The transfer of orbital angular momentum, or the transfer of high-circularity stars from the satellite to the disk may thus contribute to the early formation of the thin disk (Jiang et al., 2025). Similar variations were identified by Chen et al. (2021a) and Chen et al. (2021b) by applying a set of successively refined regressors to the galaxies in the IllustrisTNG and EAGLE simulations, where nuanced factors other than halo assembly are inferred to affect star formation in galaxies.
@@
Figure 4: The quadrant diagram showing the evolution of galaxies in the $f_{\lambda}$-$\gamma$ plane. In each panel, dots show the snapshots of all galaxies in our sample, color-coded by a property labeled in the color bar. $f_{\text{gas+star}}$ in (b) are evaluated within the inner $1\,{\rm{kpc}}$ of galaxies. The growth rate of stellar bulge ($\dot{M}_{\rm bulge}$, in $\rm M_{\odot}yr^{-1}$) shown in (h) is smoothed by the LOESS method (Cleveland & Devlin, 1988), using the Python package (Cappellari et al., 2013). White path linking the dots with black edges shows the evolution of m12f. Two vertical lines show $\gamma=3/8$ and $\gamma=0$, respectively, which define the transitional regime of halo assembly. One horizontal line shows $f_{\lambda}=1$, above which self-gravity of gas is expected to prevent the formation of a dynamically cold disk. The $f_{\lambda}$-$\gamma$ plane is partitioned into four quadrants, Q1–Q4, by these lines, within which galaxies are expected to have distinct dynamical states. The FIRE-2 galaxies follow a general evolutionary trajectory in this plane, from Q1 at high $z$ to Q3 at low $z$. See §4.1 for details.
@@ 4 Physical processes driving the two-phase growth of galaxies
The first factor is the assembly rate of the host halo, as quantified by $\gamma$. A halo in its fast phase has a nearly constant concentration of $c_{\rm v}\approx 4$ (Zhao et al., 2003, 2009), reflecting a self-similar growth without changing the shape of the profile. Such an evolution of halo structure relies on a sufficient exchange of energy between the newly accreted matter and the existing matter in the halo. This is feasible only if the accretion is fast enough so that the perturbation of the gravitational potential is frequent and strong. The fast accretion of dark matter also has important implications for baryonic processes: fast accumulation of gas is expected to be associated with fast accretion of dark matter. This, together with the effective radiative cooling expected before halos reach a threshold of $M_{\rm v}\approx 10^{12}{\rm M_{\odot}}$ (see the cooling diagram in, e.g., fig. 2 of MCW24), can cause large amounts of cold gas to fall almost freely into the halo center. Rapid inflow of gas not only sustains a high gas fraction that leads to the formation of a self-gravitating gas cloud (SGC) before angular-momentum support can be established, but also strongly perturbs the SGC and generates supersonic turbulence on the galactic scale. Density fluctuations are seeded by the gravitational perturbations and amplified by shocks in the turbulent SGC, causing the formation of sub-clouds that are dense enough to be free of cloud-cloud collision and drag force from surrounding gas. The almost ballistic motion of sub-clouds makes the mixing of angular momentum inefficient, prevents the formation of a dynamically cold disk and preserves the hot gas dynamics. Star clusters can form within the sub-clouds and inherit the hot dynamics, leading to the formation of a bulge-like stellar system. A dynamically cold gaseous and stellar disk is expected to form only after the halo assembly slows down ($\gamma\lesssim\gamma_{\rm f}$) and the gas fraction is reduced by, e.g. feedback, to a level that is comparable to the gas spin ($f_{\lambda}\equiv f_{\rm gas}/\lambda_{\rm gas}\lesssim 1$). This scenario predicts that the galaxy has hot (velocity-dispersion supported) dynamics at high $z$ and cold (rotation supported) dynamics at low $z$.
@@ 4.1 Evolutionary trajectory in the quadrant diagram
    Q2: $\gamma\lesssim\gamma_{\rm f}$ and $f_{\lambda}\gtrsim 1$. This is where unstable disks are expected to form. External perturbations from the fast halo assembly are now alleviated, and mixing of angular momentum is elevated, both promoting the formation of a self-gravitating, disk-like structure. However, local instability in the heavy disk can lead to the formation of clumps that can heat the disk dynamically (e.g. Meng & Gnedin, 2020; Romeo et al., 2020; Renaud et al., 2021; van Donkelaar et al., 2022; Romeo et al., 2023). Global instability of the disk can drive rapid gas inflow towards the galactic center to form a bar or pseudo-bulge with a certain level of dynamic hotness (e.g. Du et al., 2020; Zana et al., 2022).
@@
Figs. 4(c)–(g) show the evolution of the burstiness ($\mathcal{B}$), hotness ($\mathcal{H}$), and the fractions of the three kinematic components (bulge, thick disk, and thin disk), respectively. All these quantities exhibit obvious evolution when galaxies migrate from Q1/Q4 to Q3, revealing a general trend that galaxies evolve from a dynamically hot phase, with bulge-like morphology and highly bursty star formation, to a dynamically cold phase, with cold-disk morphology and steady star formation. All these quantities show significant correlations with $\gamma$, and make transitions at $\gamma\approx\gamma_{\rm f}=0$$3/8$. This confirms the findings in §3 that the two phases of galaxy evolution are correlated with the two phases of halo assembly. A weak trend is evident that, given $\gamma$, galaxies with higher $f_{\lambda}$ tend to have a higher hotness and bulge fraction, verifying the expected importance of $f_{\lambda}$ in determining the dynamics of galaxies. All these results support the use of the quadrant diagram to separate the phases of galaxy formation and to identify the physical processes that drive the transitions, as discussed in the following.
@@ 4.2 The origin of dynamical hotness: concentrated versus scattered star formation
The above analyses based on the quadrant diagram pinpoint two regimes for the build-up of the dynamically hot stellar system in a galaxy. The two regimes are well separated in the $f_{\lambda}$-$\gamma$ plane, and exhibit distinct patterns in the physical properties of galaxies. The later regime is at $\gamma\approx 0$$3/8$, where the fraction of cold-disk stars is about to rise (see e.g. Fig. 2a), but the SFR of the bulge component is the highest among all redshifts in the history of a galaxy (see Fig. 4h). In Fig. 5(a), we show the assembly histories of the bulge mass for all 14 galaxies in our sample. All of these assembly histories have a sigmoid-like shape, with clear transition during the transition phase of halo assembly, $t_{\gamma}^{\rm(start)}\lesssim t\lesssim t_{\gamma}^{\rm(end)}$. Bulge masses grow rapidly in this duration, and most bulge stars at $z=0$ form during this period. The strong correlation between bulge formation time ($t_{\rm bulge,1/2}$) and halo transition time ($t_{\gamma}$) shown in Fig. 3(D) is thus explained. The earlier regime is at $\gamma\gtrsim 3/8$, before the halo assembly slows down. Star formation in this regime contributes only a minor fraction of the final bulge mass at $z=0$ (see Fig. 5a), but almost all stars are formed in dynamically hot components (bulge or thick disk; see Figs. 2c and 4e, f). The two regimes thus have important implications both for observations targeting the old bulge stars at low $z$ and for deep observations targeting the ongoing bulge formation at high $z$, as we will discuss in §5.
@@
The above analyses suggest the origin of hotness in the later regime of bulge formation. Once the spatial distribution of the star-forming gas in a galaxy becomes concentrated enough so that the self-gravitation condition is satisfied, violent fragmentation of the gas naturally arises owing to various types of instabilities, such as the Jeans instability. Sub-clouds formed in this way then produce star clusters, which inherit the hot dynamics of the gas originated from the rapid inflow and the gravitational perturbations associated with the build-up of the high gas concentration. The total mass formed in this way can be very large as a result of the rich amount of gas accumulated during the fast phase of halo assembly and the deepened gravitational potential well of the host halo that resists gas loss by feedback. The formation of the stellar bulge proceeds so rapidly that a dynamically cold gaseous disk does not have enough time to form. Given the importance of high gas concentration in driving the formation of stellar bulges in this regime, we summarize the corresponding origin of the hotness as ‘concentrated star formation’. In Fig. 6(a), we show the evolution of the half-stellar-mass radii of different stellar components of m12f, in comparison to the scale radius ($r_{\rm s}$) and virial radius ($R_{\rm vir}$) of the host halo. Between $t_{\gamma}^{\rm(start)}$ and $t_{\gamma}^{\rm(end)}$, the growth of $R_{\rm vir}$ continues, but the size of the distribution of young bulge stars decreases significantly. The small size of newly formed bulge stars in comparison with the large halo size confirms the scenario of concentrated star formation inferred above from self-gravitation. At $t\gtrsim t_{\gamma}^{\rm(end)}$, the size of newly formed bulge stars drops to a few hundred parsecs, indicating that the concentrated star formation continues to take place, albeit at a very low level and within a very small region. The behavior in the evolution of the bulge size seen in m12f, is also shared by other galaxies, as shown by Fig. 6(b). In all cases, the bulge radius first increases rapidly, reaching a maximum value at $t\approx t_{\gamma}^{({\rm start})}$, and then decreases slowly with time, eventually reaching a roughly constant value at $t>t_{\gamma}^{({\rm end})}$. As shown in Fig. 6(c), when normalized by the scale radius of the host halo at the redshift when $R_{\rm 50,bulge}$ is measured, the ratio is roughly a constant of $\sim 0.3$ at $t<t_{\gamma}^{\rm(start)}$, with large fluctuations generated by the rapid change of the gravitational potential. The ratio then decreases by a factor of about three and settles around $\sim 0.1$ at $t>t_{\gamma}^{\rm(end)}$. As described in Chen et al. (2024), in the two-phase model of galaxy formation, the bulge size is related to $r_{\rm s}$ as $R_{\rm 50,bulge}\approx\alpha r_{\rm s}$, with $\alpha$ being a redshift- and mass-independent constant that is determined by the self-gravitation condition and bulge profile. Matching the prediction of the model with the observational data, Chen et al. (2024) obtained $\alpha\approx 0.04$. As one can see from Fig. 6(c), most of the simulated galaxies follow the model scaling relation after the concentrated formation of the bulge, but their values of $\alpha$ are larger by a factor of more than two. This discrepancy may partly arise from the fact that simulated bulges include all stars in the outer region where the surface brightness may be too low to be account for in the observational data. Another possibility is that the bulge to halo mass radios in simulated galaxies are larger than in real galaxies. Indeed, the stellar to halo mass ratios of FIRE-2 or FIRE-set galaxies (see Table 1; see also Hopkins et al. 2018; Schaye et al. 2025) are more than a factor of two higher than in observations (e.g. Yang et al., 2012).
@@
In the earlier regime of bulge growth, the star-forming gas is not globally self-gravitating, as seen from the low $f_{\text{gas+star}}$ in Fig. 5(b) during $t\lesssim t_{\gamma}^{\rm(start)}$. This, in combination with the high gas fractions in, and the small sizes of halos, suggests that star-forming gas must be distributed in an extended way and that the bulge size-halo size ratio is much higher than that in the later regime of bulge formation. The evolution of sizes in Fig. 6(a) clearly shows this point, where $R_{\rm 50,young\,bulge}/R_{\rm vir}$ remains as high as $0.1$ – a value much higher than 0.010.010.0150.015 observed for bulges in the local Universe (e.g. Kravtsov, 2013; Huang et al., 2017; Zanisi et al., 2020; Chen et al., 2024). These results suggest that the origin of hotness in the earlier regime is that the distribution of star-forming gas is sufficiently extended so that the random motion of gas has not been significantly dissipated until stars are born out of the gas. This is totally different from the situation where the angular momentum between gas elements is well mixed to produce a dynamically cold gaseous disk with ordered rotation before most of that gas is converted into stars. This is, however, similar to the situation in the formation of the brightest cluster galaxies (BCGs) in massive halos at low $z$, where significant fractions of stars were formed ex-situ in satellite galaxies and then accreted to the central galaxies to build their extended stellar outskirts (e.g. Bezanson et al., 2009; Huang et al., 2018; Kluge & Bender, 2023). Given the large extension of the star-forming gas and its importance in determining the hot stellar dynamics, we refer to the origin of the hotness in this regime as ‘scattered star formation’.

### v3

@@ 2.1.1 Stellar kinematics and decomposition
Here, $f_{1}$ and $f_{2}$ are the initial and final values of $f_{\rm thin}$, respectively, introduced to describe the shape of $f_{\rm thin}$ seen in Fig. 2(c) at $t<t_{1}$ and $t\geqslant t_{2}$. The change of $f_{\rm thin}$ in the transitional interval between $t_{1}$ and $t_{2}$ is described by a linear function. The fitting result for m12f is shown by the blue dashed curve in Fig. 2(c), with the duration of the transitional phase ($z\approx 0.5$11, or $t\approx 5.5$$8.5\,{\rm Gyr}$) indicated by the gray shade. To avoid ambiguity, we will denote $t_{1}$ and $t_{2}$ as $t_{\rm thin}^{\rm(start)}$ and $t_{\rm thin}^{\rm(end)}$, respectively. We use the mean value of the two,
@@ 2.2 Dark matter halos hosting the galaxies
The exact formation redshift is evaluated by linearly interpolating the redshift between the two adjacent snapshots where $M_{\rm vir}$ first crosses the half of its final mass. This time has been commonly used in the literature to characterize the assembly of dark matter halos (e.g. Gao et al., 2005; Li et al., 2008; Lyu et al., 2023, 2024), and was found to correlate with $t_{\gamma}$ (see Appendix B of MCW24). Another related formation time, which is also commonly used in the literature to characterize halo assembly, is $t_{\rm h,Vpeak}$, defined as the cosmic time when the maximum circular velocity ($V_{\rm max}$) of the halo reaches its peak along the main-branch history. This time was found to be sensitive to violent changes in the halo mass, and has been used to incorporate effects of recent major mergers (Behroozi et al., 2014).
In Table 2, we list the properties of galaxies/halos defined in this subsection and the last. Using these definitions, we obtain the characteristic times for each of the galaxies/halos in our sample, and we list them in Table 1. In general, the piecewise functions introduced in §2.1 fit well the evolution of the properties for most galaxies in our sample.
There are three exceptional cases:
  1. 1.
    m12z: This galaxy starts to evolve toward colder dynamics, but the transition has not yet been completed by $z=0$. At $z=0$, its thin disk is still subdominant ($f_{\rm thin}\approx 0.25$), and its hotness remains high ($\mathcal{H}\approx 0.5$). The ending times, $t^{\text{(end)}}_{\text{thin}}$ and $t^{\text{(end)}}_{\mathcal{H}}$, are therefore undefined for this galaxy. The star-formation burstiness of this galaxy never drops below $0.2$, and its $t_{\mathcal{B}}$ is also undefined. These properties are therefore not listed in Table 1.
  2. 2.
    m12w: This galaxy has evolved to a dynamical state colder than m12z, with $f_{\rm thin}\approx 0.65$ and $\mathcal{H}\approx 0.5$ at $z=0$. However, the transition appears to still be ongoing, with the late-time plateau value ($f_{2}$ in Eq. 2) for both $f_{\rm thin}$ and $\mathcal{H}$ determined by only a few data points at $t\gtrsim 12\,{\rm Gyr}$. The ending times, $t^{\text{(end)}}_{\text{thin}}$ and $t^{\text{(end)}}_{\mathcal{H}}$, are therefore not well constrained, and should be taken with caution for this galaxy.
  3. 3.
    m12r: The host halo of this galaxy has recent major mergers at $z\lesssim 0.3$. This introduces significant late-time jumps in the history of $V_{\rm vir}$ (and thus $\gamma$), causing the halo assembly to deviate from the typical two-phase pattern. This exception is also visible in the assembly-environment distribution (Fig. 11), where the halo of m12r appears to be an outlier.
Given the special status of these galaxies, we exclude them from the correlation analysis presented in §3. The statistics shown there should therefore only be interpreted as a result for the subset of galaxies whose halo assembly follows a typical two-phase pattern and whose galaxy dynamics has transitioned from hot to cold by $z=0$. In §4, we show that the transition of galaxy dynamics depends on other factors in addition to halo assembly, and a more detailed discussion is presented on how these three galaxies can fit in the co-evolution scenario.
@@ 3 Co-evolution pattern between galaxies and halos
Figure 3: Correlation between halo transition time ($t_{\gamma}$) and galaxy transition times. Each panel uses one definition of galaxy transition time: (A), $t_{\rm thin}$; (B), $t_{\rm\scriptscriptstyle\mathcal{H}}$; (C), $t_{\rm\scriptscriptstyle\mathcal{B}}$; (D), $t_{\rm bulge,1/2}$. Each marker represents one galaxy in our sample, with error bar spanning from the start to the end of the transition interval (for $_{\gamma}, tt$$_{\rm thin}$ and $t_{\rm\scriptscriptstyle\mathcal{H}}$) or from 20% to 80% formation time of bulge mass (for $t_{\rm bulge,1/2}$). The ending epochs of the transitions of $f_{\rm thin}$ and $\mathcal{H}$ are undefined for m12z, for which we only show the starting epochs of the transitions with arrows pointing upward. For each pair of variables, the Spearman correlation coefficient ($\rho$) is indicated in the corresponding panel (see also Table 3 for a full list of $\rho$ and $p$-values), and black line shows a linear fitting. Markers with grey edges represent the exceptional galaxies (m12w, m12r and m12z) excluded from correlation analysis and linear fitting. This figure demonstrates a co-evolution pattern between halos and galaxies. See §3 for details.
Figure 4: Durations of transitional phases for individual galaxies. Here the the duration, $\Delta T_{\rm dyn}$, is normalized by the dynamical timescale of the host halo (see Eq. 18), and the transitional phase is defined by the evolution of (a), $f_{\rm thin}$; (b), $\mathcal{H}$. Markers are the same as in Fig. 3. Grey shading indicates the range $1\leqslant\Delta T_{\rm dyn}\leqslant 3$. See §3 for details.
@@
In addition to the correlations between galaxy properties and halo assembly, there are significant variations among individual galaxies, indicating that processes other than halo accretion may also play important roles in determining the transitions in galaxy properties. For example, the galaxy m12m has a $t_{\rm thin}$ about $2\,{\rm Gyr}$ earlier than that expected from the linear fit. A careful examination reveals that m12m has a massive satellite whose orbital motion aligns with the disk rotation of m12m at $z\approx 3$. The transfer of orbital angular momentum, or the transfer of high-circularity stars from the satellite to the disk may thus contribute to the early formation of the thin disk (Jiang et al., 2025). Similar variations were identified by Chen et al. (2021b) and Chen et al. (2021a) by applying a set of successively refined regressors to the galaxies in the IllustrisTNG and EAGLE simulations, where nuanced factors other than halo assembly are inferred to affect star formation in galaxies.
@@
Figure 5: The quadrant diagram showing the evolution of galaxies in the $f_{\lambda}$-$\gamma$ plane. In each panel, dots show the snapshots of all galaxies in our sample, color-coded by a property labeled in the color bar. $f_{\text{gas+star}}$ in (b) are evaluated within the inner $1\,{\rm{kpc}}$ of galaxies. The growth rate of stellar bulge ($\dot{M}_{\rm bulge}$, in $\rm M_{\odot}yr^{-1}$) shown in (h) is smoothed by the LOESS method (Cleveland and Devlin, 1988), using the Python package (Cappellari et al., 2013). White path linking the dots with black edges shows the evolution of m12f. Two vertical lines show $\gamma=3/8$ and $\gamma=0$, respectively, which define the transitional regime of halo assembly. One horizontal line shows $f_{\lambda}=1$, above which self-gravity of gas is expected to prevent the formation of a dynamically cold disk. The $f_{\lambda}$-$\gamma$ plane is partitioned into four quadrants, Q1–Q4, by these lines, within which galaxies are expected to have distinct dynamical states. The FIRE-2 galaxies follow a general evolutionary trajectory in this plane, from Q1 at high $z$ to Q3 at low $z$. See §4.1 for details.
@@ 4 Physical processes driving the two-phase growth of galaxies
The first factor is the assembly rate of the host halo, as quantified by $\gamma$. A halo in its fast phase has a nearly constant concentration of $c_{\rm v}\approx 4$ (Zhao et al., 2003, 2009), reflecting a self-similar growth without changing the shape of the profile. Such an evolution of halo structure relies on a sufficient exchange of energy between the newly accreted matter and the existing matter in the halo. This is feasible only if the accretion is fast enough so that the perturbation of the gravitational potential is frequent and strong. The fast accretion of dark matter also has important implications for baryonic processes: fast accumulation of gas is expected to be associated with fast accretion of dark matter. This, together with the effective radiative cooling expected before halos reach a threshold of $M_{\rm v}\approx 10^{12}{\rm M_{\odot}}$ (see the cooling diagram in, e.g., fig. 2 of MCW24), can cause large amounts of cold gas to fall almost freely into the halo center. Rapid inflow of gas sustains a high gas fraction that leads to the formation of a self-gravitating gas cloud (SGC) before angular-momentum support can be established. Rapid change of gravitational potential is complemented by gas accretion and feedback processes in disturbing the gas and driving supersonic turbulence on galactic scale (as discussed in §1). Density fluctuations in the SGC are thus seeded and amplified, causing the formation of sub-clouds that are dense enough to be free of cloud-cloud collision and drag force from surrounding gas. The almost ballistic motion of sub-clouds makes the mixing of angular momentum inefficient, prevents the formation of a dynamically cold disk and preserves the hot gas dynamics. Star clusters can form within the sub-clouds and inherit the hot dynamics, leading to the formation of a bulge-like stellar system. A dynamically cold gaseous and stellar disk is expected to form only after the halo assembly slows down ($\gamma\lesssim\gamma_{\rm f}$) and the gas fraction is reduced by, e.g. feedback, to a level that is comparable to the gas spin ($f_{\lambda}\equiv f_{\rm gas}/\lambda_{\rm gas}\lesssim 1$). This scenario predicts that the galaxy has hot (velocity-dispersion supported) dynamics at high $z$ and cold (rotation supported) dynamics at low $z$.
@@ 4.1 Evolutionary trajectory in the quadrant diagram
    Q2: $\gamma\lesssim\gamma_{\rm f}$ and $f_{\lambda}\gtrsim 1$. This is where unstable disks are expected to form. External perturbations from the fast halo assembly are now alleviated, and mixing of angular momentum is elevated, both promoting the formation of a self-gravitating, disk-like structure. However, local instability in the heavy disk can lead to the formation of clumps that can heat the disk dynamically (e.g. Meng and Gnedin, 2020; Romeo et al., 2020; Renaud et al., 2021; van Donkelaar et al., 2022; Romeo et al., 2023). Global instability of the disk can drive rapid gas inflow towards the galactic center to form a bar or pseudo-bulge with a certain level of dynamic hotness (e.g. Du et al., 2020; Zana et al., 2022).
@@
Figs. 5(c)–(g) show the evolution of the burstiness ($\mathcal{B}$), hotness ($\mathcal{H}$), and the fractions of the three kinematic components (bulge, thick disk, and thin disk), respectively. All these quantities exhibit obvious evolution when galaxies migrate from Q1/Q4 to Q3, revealing a general trend that galaxies evolve from a dynamically hot phase, with bulge-like morphology and highly bursty star formation, to a dynamically cold phase, with cold-disk morphology and steady star formation. All these quantities show significant correlations with $\gamma$, and make transitions at $\gamma\approx\gamma_{\rm f}=0$$3/8$. This confirms the findings in §3 that the two phases of galaxy evolution are correlated with the two phases of halo assembly. A weak trend is evident that, given $\gamma$, galaxies with higher $f_{\lambda}$ tend to have a higher hotness and bulge fraction, verifying the expected importance of $f_{\lambda}$ in determining the dynamics of galaxies. All these results support the use of the quadrant diagram to separate the phases of galaxy formation and to identify the physical processes underlying the transitions, as discussed in the following.
@@ 4.2 The origin of dynamical hotness: concentrated versus scattered star formation
The above analyses based on the quadrant diagram pinpoint two regimes for the build-up of the dynamically hot stellar system in a galaxy. The two regimes are well separated in the $f_{\lambda}$-$\gamma$ plane, and exhibit distinct patterns in the physical properties of galaxies. The later regime is at $\gamma\approx 0$$3/8$, where the fraction of cold-disk stars is about to rise (see e.g. Fig. 2a), but the SFR of the bulge component is the highest among all redshifts in the history of a galaxy (see Fig. 5h). In Fig. 6(a), we show the assembly histories of the bulge mass for all 14 galaxies in our sample. All of these assembly histories have a sigmoid-like shape, with clear transition during the transition phase of halo assembly, $t_{\gamma}^{\rm(start)}\lesssim t\lesssim t_{\gamma}^{\rm(end)}$. Bulge masses grow rapidly in this duration, and most bulge stars at $z=0$ form during this period. The large correlation coefficient, $\rho=0.918$, between bulge formation time ($t_{\rm bulge,1/2}$) and halo transition time ($t_{\gamma}$) shown in Fig. 3(D) is thus explained. The earlier regime is at $\gamma\gtrsim 3/8$, before the halo assembly slows down. Star formation in this regime contributes only a minor fraction of the final bulge mass at $z=0$ (see Fig. 6a), but almost all stars are formed in dynamically hot components (bulge or thick disk; see Figs. 2c and 5e, f). The two regimes thus have important implications both for observations targeting the old bulge stars at low $z$ and for deep observations targeting the ongoing bulge formation at high $z$, as we will discuss in §5.
@@
The above analyses suggest the origin of hotness in the later regime of bulge formation. Once the spatial distribution of the star-forming gas in a galaxy becomes concentrated enough so that the self-gravitation condition is satisfied, violent fragmentation of the gas naturally arises owing to various types of instabilities, such as the Jeans instability. Sub-clouds formed in this way then produce star clusters, which inherit the hot dynamics of the gas originated from the perturbations and turbulence associated with the build-up of the high gas concentration. The total mass formed in this way can be very large as a result of the rich amount of gas accumulated during the fast phase of halo assembly and the deepened gravitational potential well of the host halo that may resist gas loss by feedback. The formation of the stellar bulge proceeds so rapidly that a dynamically cold gaseous disk does not have enough time to form. Given the importance of high gas concentration in driving the formation of stellar bulges in this regime, we summarize the corresponding origin of the hotness as ‘concentrated star formation’. In Fig. 7(a), we show the evolution of the half-stellar-mass radii of different stellar components of m12f, in comparison to the scale radius ($r_{\rm s}$) and virial radius ($R_{\rm vir}$) of the host halo. Between $t_{\gamma}^{\rm(start)}$ and $t_{\gamma}^{\rm(end)}$, the growth of $R_{\rm vir}$ continues, but the size of the distribution of young bulge stars decreases significantly. The small size of newly formed bulge stars in comparison with the large halo size confirms the scenario of concentrated star formation inferred above from self-gravitation. At $t\gtrsim t_{\gamma}^{\rm(end)}$, the size of newly formed bulge stars drops to a few hundred parsecs, indicating that the concentrated star formation continues to take place, albeit at a very low level and within a very small region. The behavior in the evolution of the bulge size seen in m12f, is also shared by other galaxies, as shown by Fig. 7(b). In all cases, the bulge radius first increases rapidly, reaching a maximum value at $t\approx t_{\gamma}^{({\rm start})}$, and then decreases slowly with time, eventually reaching a roughly constant value at $t>t_{\gamma}^{({\rm end})}$. As shown in Fig. 7(c), when normalized by the scale radius of the host halo at the redshift when $R_{\rm 50,bulge}$ is measured, the ratio is roughly a constant of $\sim 0.3$ at $t<t_{\gamma}^{\rm(start)}$, with large fluctuations generated by the rapid change of the gravitational potential. The ratio then decreases by a factor of about three and settles around $\sim 0.1$ at $t>t_{\gamma}^{\rm(end)}$. As described in Chen et al. (2024), in the two-phase model of galaxy formation, the bulge size is related to $r_{\rm s}$ as $R_{\rm 50,bulge}\approx\alpha r_{\rm s}$, with $\alpha$ being a redshift- and mass-independent constant that is determined by the self-gravitation condition and bulge profile. Matching the prediction of the model with the observational data, Chen et al. (2024) obtained $\alpha\approx 0.04$. As one can see from Fig. 7(c), most of the simulated galaxies follow the model scaling relation after the concentrated formation of the bulge, but their values of $\alpha$ are larger by a factor of more than two. This discrepancy may partly arise from the fact that simulated bulges include all stars in the outer region where the surface brightness may be too low to be account for in the observational data. Another possibility is that the bulge to halo mass radios in simulated galaxies are larger than in real galaxies. Indeed, the stellar to halo mass ratios of FIRE-2 or FIRE-set galaxies (see Table 1; see also Hopkins et al. 2018b; Schaye et al. 2025) are more than a factor of two higher than in observations (e.g. Yang et al., 2012).
@@
In the earlier regime of bulge growth, the star-forming gas is not globally self-gravitating, as seen from the low $f_{\text{gas+star}}$ in Fig. 6(b) during $t\lesssim t_{\gamma}^{\rm(start)}$. This, in combination with the high gas fractions in, and the small sizes of halos, suggests that star-forming gas must be distributed in an extended way and that the bulge size-halo size ratio is much higher than that in the later regime of bulge formation. The evolution of sizes in Fig. 7(a) shows this point, where $R_{\rm 50,young\,bulge}/R_{\rm vir}$ remains as high as $0.1$ – a value much higher than 0.010.010.0150.015 observed for bulges in the local Universe (e.g. Kravtsov, 2013; Huang et al., 2017; Zanisi et al., 2020; Chen et al., 2024). These results suggest that the origin of hotness in the earlier regime is that the distribution of star-forming gas is sufficiently extended so that the random motion of gas has not been significantly dissipated until stars are born out of the gas. This is totally different from the situation where the angular momentum between gas elements is well mixed to produce a dynamically cold gaseous disk with ordered rotation before most of that gas is converted into stars. This is, however, similar to the situation in the formation of the brightest cluster galaxies (BCGs) in massive halos at low $z$, where significant fractions of stars were formed ex-situ in satellite galaxies and then accreted to the central galaxies to build their extended stellar outskirts (e.g. Bezanson et al., 2009; Huang et al., 2018b; Kluge and Bender, 2023). Given the large extension of the star-forming gas and its importance in determining the hot stellar dynamics, we refer to the origin of the hotness in this regime as ‘scattered star formation’.